During the typhoon the temperature at the middle of the sea was recorded at 20 ºC . What is the temperature reading in ºR?

a. 293.16 ºR b. 440.16 ºR c. 480.00 ºR d. 528.00 ºR



1
Expert's answer
2021-04-07T10:08:26-0400
2095+491.67=528 R.20^{\circ}\cdot\dfrac{9}{5}+491.67=528\ ^{\circ}R.

Answer:

d. 528.00 ºR